OpenStudy (keana):

Given the following perfect square trinomial, fill in the missing term. 4x2 + ___x + 49

OpenStudy (keana):

4x^2

OpenStudy (anonymous):

Do you understand how to complete the square? If it was a^2 + bx + c (b/2)^2 = c So that means the b = sqrt(c) * 2 b = sqrt(49) * 2 b = what?

OpenStudy (anonymous):

You find the square root of the third term or c^2 and double it. √49 = 7 * 2 = ?

OpenStudy (keana):

is it 28 ?

OpenStudy (anonymous):

No. What's 7 * 2?

OpenStudy (keana):

that shows up as a square so are you asking what 7 & 2 is multiplied ?

OpenStudy (anonymous):

actually you're right.

OpenStudy (keana):

yay lol

OpenStudy (anonymous):

It's 28.